OpenStudy (anonymous):

Two fewer than a number doubled is the same as the number decreased by 38. Find the number. If n is "the number," which equation could be used to solve for the number?

OpenStudy (dmndlife24):

If we use "n" for the number... 2(n - 2) = n - 38 distribute... 2n - 4 = n - 38 add 4 and subtract n from both sides... n = -38 + 4 -38 + 4 = ?